OpenStudy (anonymous):

Algebra 1 Help Divide 16x^6-12x^4+4x^2 by 4x^2.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Combining like terms?

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/equation/ go here and type the equation the way you need to

OpenStudy (skullpatrol):

Combining like terms is used for adding polynomials.

OpenStudy (anonymous):

Then I have no idea.

OpenStudy (skullpatrol):

Here you want to divide each term of the polynomial by the monomial and add the results.

OpenStudy (anonymous):

Oh, okay. Thank you!

OpenStudy (skullpatrol):

$$(16x^6-12x^4+4x^2 ) \div 4x^2.$$

OpenStudy (anonymous):

*solving* so hold please:D

OpenStudy (skullpatrol):

$$(16x^6-12x^4+4x^2 ) \div 4x^2 = (16x^6-12x^4+4x^2 )* \frac{1}{4x^2} $$

OpenStudy (skullpatrol):

By definition $$a\div b= \frac{a}{b}=a*\frac{1}{b}$$$$b\neq0$$

OpenStudy (skullpatrol):

Also the distributive property allows you to write $$(b+c)a=ba+ca$$